HUBBLE SPACE TELESCOPE DAILY REPORT #2214 PERIOD COVERED: 0000Z (UTC) 09/22/98 - 0000Z (UTC) 09/23/98 Daily Status Report as of 266/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. During the final iteration of this proposal, the acquisition failure described in yesterday's report and in HSTAR 6690 occurred, affecting eight observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed STIS/CCD/MA2 8070 (Throughput of NUV filters) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of HS2027+065 in order to measure the throughput of the near ultraviolet filters. The purpose is to measure the absolute sensitivity of the NUVMAMA imaging modes with the CN182, CN270, MGII, and CIII filters. A wave length calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6827 (The Enigmatic Soft Gamma-Ray Repeaters) The WF/PC-2 was to be used to make observations of object N49-052601, the optical/ultraviolet counterpart to a enigmatic soft gamma-ray Repeater. This is the most enigmatic of a class of sources, which, for reasons unknown, repeatedly show high-energy bursts. During this proposal, the acquisition failure described in yesterday's report and in HSTAR 6690 occurred, affecting twenty-six observations. 1.5 Completed Three Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. During the first iteration of this proposal, the acquisition failure described in yesterday's report and in HSTAR 6690 occurred, affecting seven observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.6 Completed Three Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.7 Completed Five Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.9 Completed Six Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.10 Completed Four Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat fielding calibration observations in the spectroscopic mode. The observations were completed as planned, and no anomalies were reported. 1.12 Completed NIC/2 7897 (Probing Planetary Formation Around Main-Sequence Stars: A Snapshot Survey) The Near Infrared Camera (Camera Two) was used to make snapshort survey images of the main sequence star HR333 probing for planetary formation. The IR excess main-sequence stars host cool circumstellar dust disks, which are most likely replenished by the debris of unseen planetesimals. The observations were completed as scheduled, and no problems were noted. 1.13 Completed FGS 7597 (Long Term Stability of Transfer Scan Mode in FGS-1R) The Fine Guidance Sensors (FGS-1R) was used to make observations of the standard star UPGREN69 in order to study the long term stability of the Transfer Scan Mode. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/2 7834 (A Search for Giant Planets Around Very Young Nearby Late-type Dwarfs) The Near Infrared Camera (Camera Two) was used to make observations of the very young, late-type, dwarf star HD36705 as part of a search for large planets. Very young late-type dwarfs at distances closer than 30 pc are excellent sites to provide a direct detection of extrasolar planets. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/1 7870 (NICMOS Observations of FSC10214+4724: Mapping a Redshift 2.3 Quasar With 100 Parsec Resolution) The Near Infrared Camera (Camera One) was used to make mapping observations of quasar FSC10214+4 (Which has a redshift of 2.3). Our existing WF/PC-2 data provides compelling evidence that th1s luminous IRAS source is a gravitationally lensed quasar. The observations were completed as planned, and no anomalies were reported. 1.16 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared two times. As documented in HSTAR 6691, there was a NICMOS single bit error at 265/224956Z. The internal log was dumped and the ground system limit was modified. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH